OpenStudy (anonymous):

I'm having a difficult time finding inverse functions for polynomials. I know I'm close. Find the inverse function for: h(t)=-5t^2+10t+2 Rewrite: h=-5t^2+10t+2 Swap variables: t=-5h^2+10h+2 Solve for h: 5h^2-10h=2-t 5(h^2-2h)=2-t h^2-2h=(2-t)/5 h(h-2)=(2-t)/5 And this is where I'm stuck.

OpenStudy (radar):

It looks like you were going the right direction. Lets go to the next to last line of your post.:\[h ^{2}-2h =(2-t)/5\] Now I would use the "completing the square" method to solve this;\[h ^{2}-2h +1 = (2-5)/5 + 5/5 = (7-t)/5\] Notice that I added a "1" to the left hand side to make that a perfect square, but I had to also add a 1 or 5/5 to the right hand side to maintain the equality. I further simplified the right hand side. You now must take the square root of both sides. This gives you:\[h-1 =\sqrt{(7-t)/5}\] I am sure you can now take this from here and solve.

OpenStudy (anonymous):

Awesome... thanks!

OpenStudy (radar):

You're welcome, good luck with your studies.